1/x+1/y=5
2/x+3/y=13

Answer by fractalier(6550)

From
1/x + 1/y = 5
2/x + 3/y = 13
double the first one and subtract it from the second and get
2/x + 3/y = 13
-(2/x + 2/y = 10)
-----------------
1/y = 3
so that y = 1/3 and then
1/x + 3 = 5
1/x = 2
x = 1/2
(1/2 , 1/3)